If y=24 and x=784421 then what will be the answer of y-x+x×y÷xy
 

Grade:7

2 Answers

Arun
4 years ago
Dear student
 
if x = 784421 and y = 24
 
then y – x + xy /xy
 
y – x + x/x = y – x + 1 = 24 – 784421 + 1 = -784396
 
Hope it helps
 
In case ofnany difficulty, Please feel free to ask.
Vikas TU
4 years ago
Dear student 
The xy /xy both term will cancel out each other. 
So the ans of this question would be y-x 
Hope this will help 
Good Luck